solve for x and y for beer lambert equation
0.558= 0.0942(x)+ 0.0243(y) 0.309=0.0076(x)+0.047(y)
Answer #1

0.558 = 0.0942000) + 0.0243 19 ) + 0 0.309 = 0.0076 (X) + 0.047(4) es 6 multibly o equation by 0.041 we get 0.026226 = 0.0044please rate